Chapter 4, Section 4.6, Question 24 1 0 0 The matrix P-0 5 4 is the transition matrix from what basis 8- V1 V2, V3 to the basis (1,1,1),(1,1,0),(1,0,0)) for R*? 0 3 3 Click here to enter or edit your answer V1 Click here to enter or edit your answer 12 -(?? D Click here to enter or edit your answer V3 Click if you would like to Show Work for this questinen Show Work
Answer #1

Let v1 = (x1,x2,x3)T, v2 =( y1,y2,y3)T and v3 =( z1,z2,z3)T. Also let A =

1

1

1

x1

y1

z1

1

1

0

x2

y2

z2

1

0

0

x3

y3

z3

The RREF of A (on using the RREF calculator at https://www.emathhelp.net/calculators/linear-algebra/reduced-row-echelon-form-rref-caclulator/) is

1

0

0

x3

y3

z3

0

1

0

x2- x3

y2- y3

z2- z3

0

0

1

x1-x2

y1-y2

z1-z2

Hence

x3

y3

z3

x2- x3

y2- y3

z2- z3

x1-x2

y1-y2

z1-z2

=

1

0

0

0

5

4

0

3

3

Then, we have x3 = 1, x2-x3 = 0, x1-x2=0 so that x2 = 1,and x1 = 1.

Also, y3 = 0, y2- y3= 5 and y1-y2 =3 so that y2 =5 and y1 =8.

and z3 = 0, z2- z3 = 4 and z1-z2 =3 so that z2 =4 and z1 =7.

Thus, v1 = (1,1,1)T, v2 =(8,3,0)T and v3 =(7,4,0)T .